OpenStudy (anonymous):

Give an equation of the line through the point (2, 5) with slope –3/4.

OpenStudy (mathstudent55):

The equation of a line in point-slope form is y = mx + b where m = slope, and b = y-intercept.

OpenStudy (yanasidlinskiy):

Ok. Wow. Read it wrong. Sorry about that.

OpenStudy (mathstudent55):

You are given the slope, so rewrite the equation I gave you above, and substitute the value of the slope for m. Do that and show what you get.

OpenStudy (mathstudent55):

Then substitute the x- and y-coordinates of the point you are given for x and y, and calculate b. Finally, rewrite the equation using the given slope and the y-intercept you calculated.

OpenStudy (anonymous):

OK so y=(-3/4)x+5 ?

OpenStudy (mathstudent55):

You start by just substituting m with the actual slope you were given \(y = -\dfrac{3}{4}x + b\) Now to find b, we use the x- and y-coordinates of the given point and we solve for b. We use the equation with slope -3/4, and we plug in 2 for x and 5 for y. \(5 = -\dfrac{3}{4} \times 2 + b\) \(5 = -\dfrac{3}{2} + b\) \(5 +\dfrac{3}{2} = b\) \(\dfrac{10}{2} +\dfrac{3}{2} = b\) \(\dfrac{13}{2} = b\) Now that we know the value of b, we plug it back into our equation with the known value of the slope. \(y = -\dfrac{3}{4}x +\dfrac{13}{2} \)

OpenStudy (anonymous):

Ohhh thanks!

OpenStudy (mathstudent55):

You're welcome.